the number of items on two grocery lists differ by 7. the total number of items is 33. how many items are on each list?
Answer by checkley79(3341)

X-Y=7
X+Y=33 ADD.
----------------
2X=40
X=40/2
X=20 ITEMS ON THE LARGEST LIST.
20+Y=33
Y=33-20
Y=13 ANS. FOR THE SHORTEST LIST.
PROOF:
20-13=7
7=7